Show all work. Just the answer, without supporting work, will receive no credit.

The heights of pecan trees are normally distributed with a mean of 10 feet and a standard deviation of 2
feet.

1. What is the probability that a randomly selected pecan is between 8 and 12 feet tall?
2. Find the 80th percentile of the pecan tree height distribution.
3. If a random sample of 64 pecan trees is selected, what is the standard deviation of the sample mean?

Answer by ewatrrr(24785)

mean of 10 feet and a standard deviation of 2
1. P (8 > x < 12) = P(z ≤ 2/2) - P(z ≤ -2/2 ) = .8413 - .1587
2 z = invNorm(.80)= .8416 = (X-10)/2, 2(.8416) + 10 = X
3. s = 2/Sqrt(64) = .25